# Parameter estimation of gravitational-wave signals with frequency-dependent antenna responses and higher-modes

### v1

@@ 4.3 Inference using a network of dectectors
Figure 4: The antenna response as a function of frequency for GW1 and GW2 at the three detectors is analyzed. For CE and CE20, the response varies at low frequencies due to Earth’s rotation and at high frequencies due to detector-size effects. In the case of A1, the signal enters the band at 20 Hz, resulting in no frequency dependence at low frequencies. Additionally, due to the detector’s size, there is no effect at high frequencies.
  • Intrinsic parameters: The network of CE and A1 performs better than only one CE, justifying operating one present-generation detector at design sensitivity with a next-generation detector. The CEA1 network is almost at par with the CECE20 and CECE20A1 networks for GW1 and GW2 as seen in 6(c) and 6(d). This is because both these events are of sufficiently high SNR and the phase can be measured accurately in A1. This is not expected to be true at lower SNRs. Also, we do not see much additional benefit from including higher modes of radiation for intrinsic parameters.
  • Extrinsic parameters: Unlike the case for intrinsic parameters, a network of CE and CE20 is much better than a network of CE and A1 and is comparable to a network of CECE20A1. This is because for CE and CE20 Earth’s rotation has a significant effect on the amplitude (see figure 4), which is reflected in the recovered posteriors. Another interesting fact is that we see two peaks separated by an angle of $\pi/2$ for polarization angle in GW1 for a network of CECE20 and CECE20ET instead of recovering the prior as in the case of using 1 CE or in the CEA1. This is because the CECE20 and CECE20A1 are sensitive to terms of order $\mathcal{O}(\theta_{\rm JN}^{2})$. At this order, the amplitude is neither independent of $\psi$ nor is there a $\psi$, $\phi_{c}$ degeneracy. However the waveform is invariant under the transformation $\psi\hookrightarrow\psi+\frac{\pi}{2}$, which is a result of $|F_{+}|,|F_{\times}|$ and $\frac{F_{\times}}{F_{+}}$ being invariant under the same transformation.
@@
Figure 5: Inferred posteriors of extrinsic parameters of GW1 in networks comprising of:- a CE and A1 (blue); CE and CE20 (orange); and CE, CE20 and A1 (green).
@@ 4.4 Localization
Long signals, detector size effects, higher modes, and multiple detectors help in localization. In the case of only one CE, we see a 26% improvement in the 90% sky area and a 60% improvement in 3D volume for GW1 after adding the (3,3) modes. The gain in localization with higher modes and 1 CE is much greater for GW2 as the higher modes of radiation are amplified because the source is edge-on. We see a 50% improvement in sky area and a 75% improvement in sky volume after the addition of the (3,3) mode. The detailed values are given in table 3. However, the gain due to higher modes is completely lost as we add another detector to form a baseline.
@@
(2,2), (3,2) 234 <0.5 2668 <3 1677 <3 16235 <3
(2,2), (3,2), (3,3) 173 <0.5 1092 <2 829 <3 4091 <3
(2,2), (3,2), (3,3) & (4,4) 149 <0.5 744 <2 985 <3 2778 <3
Table 3: The 90% sky areas and 90% 3d sky volumes for GW1 and GW2 in one 40km CE and a network of CE and A1. Note the injections have all modes of radiation.
@@ 5 Conclusions
In this paper, we have developed the tools necessary to generate gravitational waveforms for next-generation detectors, incorporating effects due Earth’s rotation, detector size, and higher-order modes, all within the Bilby framework. To handle these waveforms, we implemented four new likelihood classes: one exact and three approximations.
@@
Using the IMRPhenomXPHM waveform family, we simulated gravitational wave signals for next-generation detectors. As BNS waveforms with higher-order modes are not yet available in the literature, we use BBH waveforms which essentially sets the tidal parameters to zero for these simulations. However, our framework supports the inclusion of tidal effects without additional modifications. For PE, we adopted aligned spin priors and used the computational efficient mode-by-mode relative binning likelihood to explore the ten-dimensional posterior space. The fiducial waveform was matched to the injected waveform for simplicity, recognizing that this assumption is infeasible in practical applications. However, this approach does not impact the primary goal of this work, which is to evaluate measurement accuracies in next-generation detectors rather than to develop PE algorithms for future detector operations. These PE runs, parallelized across 16 threads on a computing cluster, required approximately one day of wall-clock time. Additional computational optimizations, such as cython-based implementations of beam-pattern functions, could further enhance efficiency, as suggested in the literature [48].
@@
We conducted illustrative parameter estimation for two GW170817-like sources: one nearly face-on and the other edge-on. Injections included all modes, while PE runs were performed with and without the subdominant modes. For a single 40 km Cosmic Explorer (CE) detector, the inclusion of HMs improved parameter recovery. However, this advantage diminished in a network configuration of detectors. Regardless, HMs proved essential for accurately measuring the polarization angle and phase, avoiding degeneracies in both cases.
@@
The posteriors obtained by incorporating all modes of radiation appear fairly Gaussian and lack any multimodalities. Therefore, we anticipate that the Fisher formalism would provide an accurate approximation. However, this remains to be explored in future work.
@@
The accuracy controlling free parameters of our approximations were calibrated to ensure that errors remain smaller than the statistical uncertainties in PE for SNRs around 1000. For lower SNRs, where statistical uncertainties are naturally larger, the approximation parameters can be adjusted to reduce computational costs further. Additionally, at lower SNRs, the sampling process is easier and quicker. Our current methodology assumes the presence of a single signal in the data; extensions will be necessary to address overlapping signals expected in low-SNR regimes.
@@
We employed the displacement PSD, scaled by the detector length, to ensure that the noise PSD remains free from high-frequency corrections. While this paper primarily focuses on methods for studying CBCs in next-generation detectors, the proposed framework is adaptable to minor variations in PSDs or detector locations, which are yet to be finalized.
@@
$Data\leavevmode\nobreak\ and\leavevmode\nobreak\ code\leavevmode\nobreak\ availability:$ The codes for this project are hosted in https://git.ligo.org/pratyusava.baral/ce_hm_paper.
@@ 6 Acknowledgement
This work carries LIGO document number P2500078. The authors are thankful to Anson Chen for providing useful insights during internal LIGO review. This work was supported partially by the National Science Foundation (NSF) awards PHY-2207728 and PHY-2110576 and partially by Wisconsin Space Grant Consortium Awards RFP23_2-0 and RFP24_3-0. SM acknowledges support from JSPS Grant-in-Aid for Transformative Research Areas (A) No. 23H04891 and No. 23H04893. The authors are grateful for computational resources provided by the LIGO Laboratory and those provided by Cardiff University, and funded by an STFC grant supporting UK Involvement in the Operation of Advanced LIGO. PB is also grateful to the XG Mock Data Challenge workshop at Penn State University in May 2024.

### v2

@@ 4.3 Inference using a network of dectectors
Figure 5: The antenna response for the m=2 mode as a function of frequency for GW1 and GW2 at the three detectors is analyzed. For CE and CE20, the response varies at low frequencies due to Earth’s rotation and at high frequencies due to detector-size effects. In the case of A1, the signal enters the band at 20 Hz, resulting in no frequency dependence at low frequencies. Additionally, due to the detector’s size, there is no effect at high frequencies. The dotted lines denote the antenna response without Earth’s rotation effects.
  • Intrinsic parameters: The network of CE and A1 performs better than only one CE, justifying operating one present-generation detector at design sensitivity with a next-generation detector. The CEA1 network is almost at par with the CECE20 and CECE20A1 networks for GW1 and GW2 as seen in the bottom panels of figure 6 and 7. This is because both these events are of sufficiently high SNR and the phase can be measured accurately in A1. This is not expected to be true at lower SNRs. Also, we do not see much additional benefit from including higher modes of radiation for intrinsic parameters.
  • Extrinsic parameters: Unlike the case for intrinsic parameters, a three-detector network performs better than a two-detector network. This is because three detectors form a larger baseline, breaking degeneracy between the intrinsic parameters. Addition of higher modes does not improve 2D sky localization when we have multiple detectors. However, we see improved luminosity distance, inclination angle and polarization angle recovery (see top panel of 6 and 7). For GW1 (see top panel of 6), none of the networks are sensitive to terms of order $\mathcal{O}(\theta_{\rm JN}^{4})$ and so we recover the priors on the polarization angle for analysis with only the m=2 mode. However, at higher SNRs the networks can become sensitive to the fourth-order term, resulting in a bimodal distribution of the polarization angle. This is because the waveform is invariant under the transformation $\psi\hookrightarrow\psi+\frac{\pi}{2}$, which is a consequence of $|F_{+}|,|F_{\times}|$ and $\frac{F_{\times}}{F_{+}}$ being invariant under the same transformation. However, higher-order modes are required to choose the correct mode of $\psi$.
@@
(c) Modes: (2,2), (3,2)
(d) Modes: (2,2), (3,2), (3,3), (4,4)
Figure 6: Inferred posteriors of extrinsic parameters (top panel) and intrinsic parameters (bottom panel) of GW1 in networks comprising of:- a CE and A1 (blue); CE and CE20 (orange); and CE, CE20 and A1 (green).
@@ 4.4 Localization
Long signals, detector size effects, higher modes, and multiple detectors all contribute to localization. Ignoring effects due to Earth’s rotation and detector size results in biased PE and poor sky localization (see figure 4) as the changes in the antenna response (see figure 5) and the Doppler shift are not properly accounted for. In the case of only one CE, we observe a 42% improvement in the 90% sky area and a 60% improvement in 3D volume for GW1 (see figure 3(a)) after adding the (3,3) and (4,4) modes. For GW2, ignoring the m=3 mode and m=4 mode results in biased parameter recovery (see figure 3(b)), as the higher modes of radiation carry significant SNR. The detailed values are given in table 3. However, the gain due to higher modes is lost as we add another detector to form a baseline.
@@
(2,2), (3,2) 578 <1.5 51732 77 - - - -
(2,2), (3,2), (3,3) 313 <1.5 18697 57 902 <3 4234 <3
(2,2), (3,2), (3,3) & (4,4) 334 <1.5 20373 55 948 <3 2468 <3
Table 3: The 90% sky areas and 90% 3d sky volumes for GW1 and GW2 in one 40km CE and a network of CE and A1. For GW2, the run containing only the m=2 mode did not recover the position of the injection and hence the sky areas and the sky volumes are not reported. Note that the injections have all modes of radiation.
@@ 5 Conclusions
In this paper, we develop the tools necessary to generate GW waveforms for next-generation detectors, incorporating effects due to Earth’s rotation, detector size, and higher-order modes, all within the Bilby framework. To handle these waveforms, we implement new likelihood classes, including approximations to the exact likelihood to speed up sampling. We are the first to perform proof-of-concept PE including all these effects.
We simulate gravitational wave signals for next-generation detectors using the IMRPhenomXPHM waveform family. For demonstration purposes, we select a GW170817-like signal, which remains in band for an extended duration, making it well-suited to highlight the significance of Earth-rotation effects. We use black hole waveforms in this analysis, thus setting tidal effects to zero. The waveforms go up to 4 kHz to incorporate detector-size effects (see figure 5) while the $f_{\rm isco}$ for the system is $\sim$ 2kHz. The merger portion of the waveform ($f>f_{\rm isco}$) depends on the neutron star constituents and hence, the high-frequency behavior is unrealistic. This does not cause any biases in our inference, since we are using the same family to generate simulated signals and perform PE. In principle, however, our framework supports the use of BNS waveforms such as TEOBResumS [53], SEOBNRv5HM [54], or NRHybSur3dq8Tidal [55] without modification, aside from the increased computational cost associated with waveform generation. Using a more realistic waveform to perform PE is left for the future.
@@
For PE, we adopted aligned spin priors and used the computationally efficient mode-by-mode relative binning likelihood to explore the ten-dimensional posterior space. The injected waveform is taken to be the fiducial waveform for simplicity, recognizing that this assumption is infeasible in practical applications. However, this approach does not impact the primary goal of this work, which is to evaluate measurement accuracies in next-generation detectors rather than to develop PE algorithms for future detectors. These PE runs, parallelized across 4 threads on a computing cluster, required approximately one day of wall-clock time. Additional computational optimizations, such as cython-based implementations of beam-pattern functions, could further enhance efficiency, as suggested in the literature [48].
Baker et al. [26] recently investigated the use of reduced order quadrature (ROQ) models to perform PE on BNS signals while accounting for Earth’s rotation and detector size effects. Their study concluded that the computational cost becomes prohibitive when analyzing signals from frequencies below 16 Hz, even when higher-order modes are ignored. However, the problem can be solved using multibanding as demonstrated in [21]. Our implementation using relative binning enables efficient likelihood evaluation starting from 5 Hz while including higher-order modes, with only negligible loss in accuracy and precision. This demonstrates the scalability and robustness of relative binning for next-generation PE.
While Baker et al. pointed out that multibanding and relative binning may not be sufficiently accurate for analyzing BNS signals with SNRs $\gtrsim 10^{3}$, we find that the log-likelihood errors introduced by these approximation methods remain below unity even for such high-SNR signals. A key difference between our study and theirs lies in how the errors are evaluated: we compute them on posterior samples obtained from parameter estimation, whereas they evaluate them on random samples drawn from the prior. Since relative binning is expected to be accurate primarily in regions where the waveform closely resembles a reference waveform and typically the posterior is concentrated, it is not necessarily accurate across the entire prior volume. Nevertheless, as demonstrated in our results, parameter estimation using relative binning accurately recovers the injected parameters, indicating that the method is sufficiently accurate in the regions that matter. For multibanding, we tune the parameters controlling the accuracy to ensure that log-likelihood errors remain below unity, while still achieving a computational speed-up of $\mathcal{O}(10^{2})$.
Our approach avoids the costly precomputation of an ROQ basis. Instead, we require only the evaluation of a fiducial waveform, whose parameters are chosen to exactly match those of the injected signal. Our method has virtually zero computational overhead. We note, however, that this idealized setup does not apply to realistic PE, where the true signal parameters are unknown. Generalizing the method to such cases, potentially by iterating on the fiducial waveform or incorporating some hybrid inference schemes using multibanding and relative binning, is to be explored in future studies. We also construct a multibanding likelihood to address the same problem, though we find it computationally expensive to sample from, albeit feasible. As demonstrated in [26], neglecting Earth’s rotation and finite-size detector effects introduces biases in parameter recovery. Nonetheless, for forecasting studies, this configuration provides an accurate and computationally efficient framework for assessing the impact of effects such as Earth’s rotation, detector size, and higher-order modes of radiation. It offers a promising foundation for developing scalable inference techniques for next-generation detectors.
@@
We conduct illustrative parameter estimation for two GW170817-like sources: one nearly face-on and the other edge-on. Injections include all modes, while PE runs are performed with and without the subdominant modes. For a single 40 km Cosmic Explorer (CE) detector, the inclusion of HMs improves parameter recovery. However, this advantage diminishes in a network comprising multiple detectors. Regardless, HMs prove essential for accurately measuring the polarization angle and phase, avoiding degeneracies in both cases.
@@
The posteriors we obtain by incorporating all modes of radiation appear fairly Gaussian and lack any multimodalities. Therefore, we anticipate that the Fisher formalism would provide an accurate approximation. However, this remains to be explored in future work.
@@
The accuracy-controlling free parameters of our approximations are calibrated to ensure that errors remain smaller than the statistical uncertainties in PE for SNRs around 1000. For lower SNRs, where statistical uncertainties are naturally larger, the approximation parameters can be adjusted to reduce computational costs further. Additionally, at lower SNRs, the sampling process is easier and quicker. Our current methodology assumes the presence of a single signal in the data; extensions will be necessary to address overlapping signals expected in low-SNR regimes.
@@
We employ the displacement PSD, scaled by the detector length, to ensure that the noise PSD remains free from high-frequency corrections. While this paper primarily focuses on methods for studying CBCs in next-generation detectors, the proposed framework is adaptable to minor variations in PSDs or detector locations, which are yet to be finalized.
@@
$Data\penalty 10000\ and\penalty 10000\ code\penalty 10000\ availability:$ The codes for this project are hosted in https://git.ligo.org/pratyusava.baral/ce_hm_paper/-/tree/same_snr?ref_type=heads. The forked version of Bilby containing all the modifications for the next generation detector is available here: https://git.ligo.org/pratyusava.baral/bilby-x-g/-/tree/freq_dep_antenna_response_HM?ref_type=heads. The data and codes are also available via Zenodo [64].
@@ 6 Acknowledgement
This work carries LIGO document number P2500078. The authors are thankful to Anson Chen for providing useful insights during the internal LIGO review. This work was supported partially by the National Science Foundation (NSF) awards PHY-2207728 and PHY-2110576 and partially by Wisconsin Space Grant Consortium Awards RFP23_2-0 and RFP24_3-0. SM acknowledges support from JSPS Grant-in-Aid for Transformative Research Areas (A) No. 23H04891 and No. 23H04893. The authors are grateful for computational resources provided by the LIGO Laboratory and those provided by Cardiff University, and funded by an STFC grant supporting UK Involvement in the Operation of Advanced LIGO. PB is also grateful to the XG Mock Data Challenge workshop at Penn State University in May 2024.